OpenStudy (mony01):

find a power series representation for the function f(x)=4+x/1-x

OpenStudy (loser66):

\[f(x) = 4 +\frac{x}{1-x}\] right?

OpenStudy (loser66):

No? ok.

OpenStudy (adamaero):

it has to be, what question has x/1

OpenStudy (mony01):

its \[f(x)=\frac{ 4+x }{ 1-x }\]

OpenStudy (loser66):

\[f(x) = (4+x )*\frac{1}{1-x}=(4+x)\sum_0^k x^k\], Dat sit

OpenStudy (loser66):

if you want, you can open the sum to get the "power series" form. Same !!

OpenStudy (loser66):

sorry, the letter above sum is infinitive, not k